OpenStudy (anonymous):

4x-5y=4 and 2x-y=8,using this information,determine the value of x-y.

OpenStudy (anonymous):

add the two equations and divide by 6.

OpenStudy (anonymous):

4x-5y=4.....(1) 2x-y=8......(2) (2)x2 4x-2y=16...(2') (2')-(1) 3y=12 y=4 sub ke (2) 2x-4=8 2x=12 x=6 x=6,y=4

OpenStudy (anonymous):

4x - 5y = 4 +2x - y = 8 ------------ 6x - 6y = 12 Dividing by 6 gives: \[\frac{6x-6y}{6} = \frac{12}{6} \iff x - y = 2\]

OpenStudy (anonymous):

thanks